OpenStudy (anonymous):

Choose the correct simplification of the expression g^5h^4/g^2h^3

OpenStudy (raden):

use the property : a^m/a^n = a^(m-n)

OpenStudy (anonymous):

\[\frac{g^5h^5}{g^2h^3} = g^{(5-2)}h^{(4-3)} = g^2h\]

OpenStudy (anonymous):

\[\frac{ g ^{5}h ^{4} }{ g ^{2}h ^{3} }\]

OpenStudy (anonymous):

what do i substitute with what @RadEn

OpenStudy (raden):